# Introducing NewCluster: the first half of the history of a high-resolution cluster simulation

### v1

@@ 3.7 Cluster Merger
The main halo in the NewCluster simulation is approaching a major merger event (mass ratio 1:3.2). The second-largest halo lies in close proximity, with its first encounter expected at $z\approx 0.774$ and final coalescence at $z\approx 0.35$, according to the DM-only pilot simulation. Figure 11 illustrates a comprehensive, multi-component view of the main and secondary halos involved in the major merger. The upper six panels display: (A) DM density, (B) stellar density, (C) gas column density, (D) gas temperature, (E) the Sunyaev–Zeldovich (SZ) signal computed using the Compton $y$-parameter, which traces the integrated pressure along the line of sight (Nelson et al., 2024), and (F) gas metallicity. The two halos are projected onto a plane containing the collision axis, providing a clear view of their relative structure and separation. Red and black dashed circles indicate the virial radii of the primary and secondary halos, respectively.
@@
The bottom right panel of Fig. 11 shows the median ram pressure experienced by galaxies along the collision axis. Notably, galaxies located between the two clusters are subject to elevated ram pressure compared to galaxies at similar distances on the far side of the merger front. This enhancement is a consequence of the ICM compression, indicating that galaxies can be affected by cluster-cluster mergers, even though they are treated as collisionless components in large-scale dynamics. Such an increased ram pressure may lead to triggered SF or AGN activity in some galaxies (Owen et al., 1999; Miller & Owen, 2003; Stroe et al., 2014; Stroe & Sobral, 2021). In this context, NewCluster provides a valuable platform for investigating the impact of large-scale mergers on galaxy evolution.
@@ 3.8 Quenching of galaxies in Clusters
Red and dead galaxies preferentially reside in high-density environments in the local Universe (Dressler, 1980; Lewis et al., 2002; Hogg et al., 2004; Kauffmann et al., 2004). In contrast, these regions likely experienced elevated star formation activity in the early Universe (Jeon et al., 2022). Observations suggest that substantial environmental quenching was already underway at high redshifts. For example, Gerke et al. (2007) reported an excess of blue galaxies in low-density regions (groups and the field) at $z\lesssim 1$. Furthermore, the emergence of a well-defined colour-magnitude relation in galaxy clusters at $z\sim 1$ (Ellis et al., 1997; Kodama et al., 1998; Stanford et al., 1998; Mei et al., 2006a, b) supports the idea of early galaxy formation in dense environments. The presence of a red sequence on the colour-magnitude plane at these redshifts is considered a hallmark of such accelerated evolution. Evidence of rapid environmental influence extends to even higher redshifts (e.g., $z>2$; Steidel et al., 2005), and JWST observations have begun tracing the formation and evolution of galaxies in overdense regions beyond $z\sim 5$ (e.g., Alberts et al., 2024; Morishita et al., 2025). These findings suggest that environmental effects may have already played a significant role in shaping galaxy populations at $z>1$.
In this context, NewCluster is particularly well suited for investigating the origin of SF quenching at high redshifts, as it extends beyond $z=1$ and contains several thousand satellite galaxies residing in high-density environments. For example, we select two halos with $M_{\rm vir}>10^{13}\,M_{\odot}$ at $z=1$. Galaxies with $M_{*}>10^{9}\,M_{\odot}$ located within three virial radii of these halos are selected and classified based on their SF activity. Specifically, galaxies are divided into star-forming ($b>0.1$) or quenched ($b<0.1$) populations, where $b=t_{\rm Universe}\times{\rm SFR}/M_{*}$ is the birth rate parameter widely used to quantify the degree of SF quenching across a broad range of redshift and stellar masses, with $t_{\rm Univserse}$ denoting the age of the Universe (Franx et al., 2008; Lotz et al., 2019; Park et al., 2022; Rhee et al., 2024). The SFR is measured over a $100\,{\rm Myr}$ timescale. At this stage, galaxies with $M_{*}<10^{9}\,M_{\odot}$ are excluded from the classification, as the majority of these low-mass galaxies exhibit negligible or no SF activity. Their rapid quenching is likely driven by their shallow gravitational potentials, which make them highly susceptible to environmental processes within clusters.
@@
The quenched galaxies (red circles) are primarily concentrated within the virial radius, whereas star-forming galaxies tend to occupy the outer regions, consistent with them being a recently-accreted population (Rhee et al., 2017). This clear separation between the two distributions suggests a scenario in which galaxies are rapidly quenched following infall into their host clusters. A similar separation between quenched and star-forming populations has been observed in clusters at the local Universe, where phase-space diagrams reveal strong correlations between the quenching status and the orbital history of galaxies (e.g., Oman & Hudson, 2016; Rhee et al., 2020; Oman et al., 2021). These findings highlight the predictive power of phase-space analysis that extends beyond the local universe and remains effective even at high redshifts (e.g., Muzzin et al., 2014; Kim et al., 2023b).
Given that the environmental effects in the clusters are already evident by $z=1$ (e.g., Gay et al., 2010), the NewCluster simulation is well positioned to support a range of follow-up studies. One important direction is to investigate the physical mechanisms driving rapid quenching, which likely account for the lack of star-forming galaxies in cluster core regions. Beyond the impact on newly-accreted satellites, the simulation also offers an opportunity to examine the evolution of the Butcher–Oemler effect (Butcher & Oemler, 1978) at high redshifts, providing insights into the pace of galaxy transformation in high-density environments. Such studies will be particularly valuable when interpreted alongside recent JWST observations tracing early galaxy populations in overdense regions.
@@ 4 Summary
This paper introduces the current status of the NewCluster simulation and outlines potential scientific investigations it enables. A cluster region is selected from a cosmological volume with a side-length of $100\,h^{-1}\,\mathrm{Mpc}$, and is subsequently re-simulated at high resolutions, incorporating a comprehensive and modern astrophysics model (see Sect. 2). In addition to the standard prescriptions for SF and feedback models from stars and super-massive black holes (SMBHs), NewCluster performs on-the-fly modeling of dust evolution and chemical enrichment for ten species (H, D, He, C, N, O, Mg, Fe, Si, and S). These capabilities allow for the investigation of previously inaccessible aspects of galaxy evolution in a cosmological context. For instance, shaping galaxy morphologies at high redshifts, informed by dust extinction from realistic dust distributions within galaxies has been studied in Byun et al. (2025, submitted). Figure 2 highlights these novel features, presenting mock images of NewCluster galaxies that closely resemble real observations.
Because it simulates one of the most complex regions of the Universe at very high resolution, NewCluster requires computationally intensive calculations. For example, the simulation reached $z=0.8$, spending more than $10^{8}$ CPU hours. These extensive calculations have been feasible thanks to the massive implementation of parallelization on the RAMSES code using hybrid parallelization of both OpenMP and MPI (Han et al., 2025). The updated RAMSES code shows efficient performance in using numerous CPU cores simultaneously, with which the NewCluster simulation is running using 7,680–15,360 threads, depending on the speed and resource situations.
@@
Advancing toward a better understanding of dust evolution through cosmic time might be one of the most important and urgent tasks in the JWST era. Because not only the chemical composition of grain species but also their geometric distribution matters, a physically motivated theoretical prediction in a cosmological context is crucial. Although many cosmological simulations have included dust evolution on-the-fly (Gjergo et al., 2018; Hou et al., 2019; Trebitsch et al., 2021; Granato et al., 2021; Lewis et al., 2023), NewCluster is one of the pioneering cosmological simulations to simultaneously track both the grain size distribution and chemical composition of dust down to redshifts $z<1$.
@@
[…]

### v2

@@ 3.7 Cluster merger
The main halo in the NewCluster simulation is approaching a major merger event (mass ratio 1:3.2). The second-largest halo lies in close proximity, with its first encounter expected at $z\approx 0.774$ and final coalescence at $z\approx 0.35$, according to the DM-only pilot simulation. Figure 11 illustrates a comprehensive, multi-component view of the main and secondary halos involved in the major merger. The upper six panels display: (A) DM density, (B) stellar density, (C) gas column density, (D) gas temperature, (E) the Sunyaev-Zeldovich (SZ) signal computed using the Compton $y$-parameter, which traces the integrated pressure along the line of sight (Nelson24), and (F) gas metallicity. The two halos are projected onto a plane containing the collision axis, providing a clear view of their relative structure and separation. Red and black dashed circles indicate the virial radii of the primary and secondary halos, respectively.
@@
The bottom right panel of Fig. 11 shows the median ram pressure experienced by galaxies along the collision axis. Notably, galaxies located between the two clusters are subject to elevated ram pressure compared to galaxies at similar distances on the far side of the merger front. This enhancement is a consequence of the ICM compression, indicating that galaxies can be affected by cluster-cluster mergers, even though they are treated as collisionless components in large-scale dynamics. Such an increased ram pressure may lead to triggered SF or AGN activity in some galaxies (Owen99; Miller03; Stroe14; Stroe21). In this context, NewCluster provides a valuable platform for investigating the impact of large-scale mergers on galaxy evolution.
@@ 3.8 Quenching of galaxies in Clusters
Red and dead galaxies preferentially reside in high-density environments in the local Universe (Dressler1980; Lewis02; Hogg04; Kauffmann04). In contrast, these regions likely experienced elevated star formation activity in the early Universe (2022ApJ...941....5J). Observations suggest that substantial environmental quenching was already underway at high redshifts. For example, Gerke07 reported an excess of blue galaxies in low-density regions (groups and the field) at $z\lesssim 1$. Furthermore, the emergence of a well-defined colour-magnitude relation in galaxy clusters at $z\sim 1$ (Ellis97; Kodama98; Stanford98; Mei06a; Mei06b) supports the idea of early galaxy formation in dense environments. The presence of a red sequence on the colour-magnitude plane at these redshifts is considered a hallmark of such accelerated evolution. Evidence of rapid environmental influence extends to even higher redshifts (e.g., $z>2$; Steidel05), and JWST observations have begun tracing the formation and evolution of galaxies in overdense regions beyond $z\sim 5$ (e.g., Alberts24; Morishita25). These findings suggest that environmental effects may have already played a significant role in shaping galaxy populations at $z>1$.
In this context, NewCluster is particularly well suited for investigating the origin of SF quenching at high redshifts, as it extends beyond $z=1$ and contains several thousand satellite galaxies residing in high-density environments. For example, we select two halos with $M_{\rm vir}>10^{13}\,M_{\odot}$ at $z=1$. Galaxies with $M_{*}>10^{9}\,M_{\odot}$ located within three virial radii of these halos are selected and classified based on their SF activity. Specifically, galaxies are divided into star-forming ($b>0.1$) or quenched ($b<0.1$) populations, where $b=t_{\rm Universe}\times{\rm SFR}/M_{*}$ is the birth rate parameter widely used to quantify the degree of SF quenching across a broad range of redshift and stellar masses, with $t_{\rm Univserse}$ denoting the age of the Universe (Franx08; Lotz19; Park22; Rhee24). The SFR is measured over a $100\,{\rm Myr}$ timescale. At this stage, galaxies with $M_{*}<10^{9}\,M_{\odot}$ are excluded from the classification, as the majority of these low-mass galaxies exhibit negligible or no SF activity. Their rapid quenching is likely driven by their shallow gravitational potentials, which make them highly susceptible to environmental processes within clusters.
@@
The quenched galaxies (red circles) are primarily concentrated within the virial radius, whereas star-forming galaxies tend to occupy the outer regions, consistent with them being a recently-accreted population (Rhee17). This clear separation between the two distributions suggests a scenario in which galaxies are rapidly quenched following infall into their host clusters. A similar separation between quenched and star-forming populations has been observed in clusters at the local Universe, where phase-space diagrams reveal strong correlations between the quenching status and the orbital history of galaxies (e.g., Oman16; Rhee20; Oman21). These findings highlight the predictive power of phase-space analysis that extends beyond the local universe and remains effective even at high redshifts (e.g., Muzzin14; Kim23).
Given that the environmental effects in the clusters are already evident by $z=1$ (e.g., Gay10), the NewCluster simulation is well positioned to support a range of follow-up studies. One important direction is to investigate the physical mechanisms driving rapid quenching, which likely account for the lack of star-forming galaxies in cluster core regions. Beyond the impact on newly-accreted satellites, the simulation also offers an opportunity to examine the evolution of the Butcher-Oemler effect (BO78) at high redshifts, providing insights into the pace of galaxy transformation in high-density environments. Such studies will be particularly valuable when interpreted alongside recent JWST observations tracing early galaxy populations in overdense regions.
@@ 4 Summary
This paper introduces the current status of the NewCluster simulation and outlines potential scientific investigations it enables. A cluster region is selected from a cosmological volume with a side-length of $100\,h^{-1}\,\mathrm{Mpc}$, and is subsequently re-simulated at high resolutions, incorporating a comprehensive and modern astrophysics model (see Sect. 2). In addition to the standard prescriptions for SF and feedback models from stars and super-massive black holes (SMBHs), NewCluster performs on-the-fly modeling of dust evolution and chemical enrichment for ten species (H, D, He, C, N, O, Mg, Fe, Si, and S). These capabilities allow for the investigation of previously inaccessible aspects of galaxy evolution in a cosmological context. For instance, shaping galaxy morphologies at high redshifts, informed by dust extinction from realistic dust distributions within galaxies has been studied in 2025ApJ...992...92B. Figure 2 highlights these novel features, presenting mock images of NewCluster galaxies that closely resemble real observations.
Because it simulates one of the most complex regions of the Universe at very high resolution, NewCluster requires computationally intensive calculations. For example, the simulation reached $z=0.8$, spending more than $10^{8}$ CPU hours. These extensive calculations have been made possible thanks to significant scalability enhancements in the RAMSES code, achieved through hybrid parallelization with OpenMP and MPI (yomp). The updated RAMSES code shows efficient performance in using numerous CPU cores simultaneously, with which the NewCluster simulation is running using 7,680–15,360 threads, depending on the speed and resource situations.
@@
Advancing toward a better understanding of dust evolution through cosmic time might be one of the most important and urgent tasks in the JWST era. Because not only the chemical composition of grain species but also their geometric distribution matters, a physically motivated theoretical prediction in a cosmological context is crucial. Although many cosmological simulations have included dust evolution on-the-fly (2018MNRAS.479.2588G; 2019MNRAS.485.1727H; 2021A&A...653A.154T; Granato21; 2023MNRAS.519.5987L), NewCluster is one of the pioneering cosmological simulations to simultaneously track both the grain size distribution and chemical composition of dust down to redshifts $z<1$.
@@
A critical shortcoming of NewCluster is that it is only one cluster. Therefore, it cannot be representative. However, it is already too challenging to get one cluster done for such resolutions. After 3 years of nearly-continuous runs, we only reached z=0.8z=0.8888As we speak, NewCluster is passing $z=0.70$., and we expect to require a similar or greater amount of computing time to reach the local universe. Now that we see the beauty and power of high-resolution data, we want to continue the march to simulate more clusters. But for that to be possible, a dramatically improved numerical approach would be necessary.